11. A bond makes annual coupon payments, is currently worth $995, and has a duration of 4.77 years. The bond's yield-to-maturity is currently 9%. If the rate on same-risk bonds suddenly changes to 8.52%, then what does duration predict the new price of the bond to be? Round your answer to the nearest penny.
Answer #1

Change in bond price/Initial bond price = -Change in yield * Duration

Change in bond price = -995*(0.0852-0.09)*4.77

Change in bond price = 22.78

New bond price = 22.78+995 = $1017.78 ~ $1018